# Entropic Risk Optimization in Discounted MDPs: Sample Complexity Bounds with a Generative Model

### v2

@@ Appendix C Bellman Optimality and Bellman Recursions
$LQ(s,a):=R(s,a)+\gamma\rho_{s,a}(\max_{a^{\prime}}Q(s^{\prime},a^{\prime}))\,.$ (36)
$LQ_{1}(s,a)$ $-LQ_{2}(s,a)=\gamma[\rho_{s,a}(\max_{a^{\prime}}Q_{1}(s^{\prime},a^{\prime}))-\rho_{s,a}(\max_{a^{\prime}}Q_{2}(s^{\prime},a^{\prime}))]$ (37)
$=\gamma[\rho_{s,a}(\max_{a^{\prime}}Q_{1}(s^{\prime},a^{\prime}))-\rho_{s,a}(\max_{a^{\prime}}Q_{1}(s^{\prime},a^{\prime})-\max_{a^{\prime}}Q_{1}(s^{\prime},a^{\prime})+\max_{a^{\prime}}Q_{2}(s^{\prime},a^{\prime}))]$ (38)
$\leq\gamma[\rho_{s,a}(\max_{a^{\prime}}Q_{1}(s^{\prime},a^{\prime}))-\rho_{s,a}(\max_{a^{\prime}}Q_{1}(s^{\prime},a^{\prime})+\max_{a^{\prime}}\{Q_{2}(s^{\prime},a^{\prime})-Q_{1}(s^{\prime},a^{\prime}))\})]$ (39)
$\leq\gamma[\rho_{s,a}(\max_{a^{\prime}}Q_{1}(s^{\prime},a^{\prime}))-\rho_{s,a}(\max_{a^{\prime}}Q_{1}(s^{\prime},a^{\prime})+\|Q_{1}-Q_{2}\|)]$ (40)
$=\gamma\|Q_{1}-Q_{2}\|\,.$ (41)
We start by showing that $L:B(\mathcal{S}\times\mathcal{A})\rightarrow B(\mathcal{S}\times\mathcal{A})$, that is it takes non-negative functions and returns non-negative functions. By normalization and monotonicity of $\rho$ we have for any non-negative $Q\in B(\mathcal{S}\times\mathcal{A})$ that
$LQ(s,a)=R(s,a)+\gamma\rho_{s,a}(\max_{a^{\prime}}Q(s,a))\geq 0+\gamma\rho_{s,a}(0)=0$ (42)
@@
Since there are only finitely many states and actions we can pick a stationary decision rule where $f^{*}(s)$ is an arbitrary element of $\arg\!\max_{a}Q(s,a)$.
Let $V$ be the function given $V(s):=Q(s,f^{*}(s))$ for all $s$. We then see that
$V(s)$ $\geq R(s,a)+\gamma\rho_{s,a}(V(s^{\prime}))$ (43)
$Q(s,a)\geq(L_{a}\circ L_{\pi_{2}}\circ\cdots\circ L_{\pi_{N}})V(s)\geq(L_{a}\circ L_{\pi_{2}}\circ\cdots\circ L_{\pi_{N}})\boldsymbol{0}(s)=J_{N}(s,a,\pi)\,$, (44)
Finally we aim to that $Q(s,a)\leq J(s,a,f^{*})$. By induction, we wish to show that $V(s)\leq J_{N}(s,f^{*}(s),f^{*})+\gamma^{N}\|V\|$ for all $N\in\mathbb{N}$. For the induction step, we start by noting that $J_{1}(s,f^{*}(s),f^{*})=R(s,f^{*}(s))$ and so
$V(s)$ $=R(s,f^{*}(s))+\gamma\rho_{s,f^{*}(s)}(V(s^{\prime}))$ (45)
$\leq R(s,f^{*}(s))+\gamma\rho_{s,f^{*}(s)}(\|V\|)$ (46)
$=R(s,f^{*}(s))+\gamma\|V\|$ (47)
$=J_{1}(s,f^{*}(s),f^{*})+\gamma\|V\|\,$, (48)
$V(s)$ $=L_{f^{*}}V(s)$ (49)
$\leq L_{f^{*}}(J_{N}(s,f^{*}(s),f^{*})+\gamma^{N}\|V\|)$ (50)
$=\big(R(\cdot,f^{*}(\cdot))+\gamma\rho_{\cdot,f^{*}(\cdot)}(J_{N}(\cdot,f^{*}(\cdot),f^{*})+\gamma^{N}\|V\|\big)(s)$ (51)
$=J_{N+1}(s,f^{*}(s),f^{*})+\gamma^{N+1}\|V\|\,$, (52)
from which taking the limit $N\rightarrow\infty$, we get that $V(s)\leq J(s,f^{*}(s),f^{*})$.
Finally, since
$Q(s,a)=L_{a}V(s)\leq L_{a}J(s,f^{*}(s),f^{*})=J(s,a,f^{*})\,$, (53)
the conclusion holds.
Since this shows that an optimal stationary policy exists, it will suffice to consider only stationary policies and one can by completely analogous arguments show that for any stationary policy $\pi$, there exists a non-negative map $Q^{\pi}\in B(\mathcal{S}\times\mathcal{A})$ such that $Q^{\pi}(s,a)=J(s,a,\pi)$ such that $Q^{\pi}$ satisfies the Bellman recursion:
$Q^{\pi}(s,a)=R(s,a)+\gamma\rho_{s,a}(Q^{\pi}(s^{\prime},\pi(s^{\prime}))\,$, (54)
We also remark that in the proof we see directly that $Q(s,a)\in[0,\frac{1}{1-\gamma}]$ for all $(s,a)$.
Appendix D Deterioration of Greedy Policy
Next we show a result that bounds the quality of a greedy policy with respect to the quality of the value-function for which the policy is greedy. The result is a generalization of [56] from the expectation to that of the ERM. Throughout, we use the notation $\rho_{s,a}(V(s^{\prime}))$ as shorthand notation for $\rho$ applied to the categorical random variable $X$ with support $\{V(s^{\prime})\}_{s^{\prime}\in\mathcal{S}}$ where $\mathbb{P}(X=V(s^{\prime}))=P(s^{\prime}|s,a)$.
Theorem 7.
Let $\widehat{V}$ be a function for which $\|V^{*}-\widehat{V}\|<\varepsilon$ and let $\pi^{G}:=\arg\!\max_{a}[R(s,a)+\gamma\rho_{s,a}(\widehat{V}(s^{\prime}))]$ be a greedy policy with respect to $\widehat{V}$. Let the value function of this greedy policy be denoted $V^{G}:=V^{\pi^{G}}$. It then holds that
$\|V^{*}-V^{G}\|\leq\frac{2\gamma}{1-\gamma}\varepsilon\,.$ (55)
@@ Appendix D Deterioration of Greedy Policy
Let $\bar{s}$ be a state such that $\|V^{*}-V^{G}\|=V^{*}(\bar{s})-V^{G}(\bar{s})$. We then consider the two actions $a^{*}:=\pi^{*}(\bar{s})$ (pick any if more) and $a^{G}:=\pi^{G}(\bar{s})$. Since $\pi^{G}$ is greedy w.r.t. $V^{G}$, we have that
$R(\bar{s},a^{*})+\gamma\rho_{\bar{s},a^{*}}(\widehat{V}(s^{\prime}))\leq R(\bar{s},a^{G})+\gamma\rho_{\widehat{s},a^{G}}(\widehat{V}(s^{\prime}))\,.$
@@
$R(\bar{s},a^{*})+\gamma\rho_{\bar{s},a^{*}}(\widehat{V}(s^{\prime}))$ $\geq R(\bar{s},a^{*})+\gamma\rho_{\bar{s},a^{*}}(V^{*}(s^{\prime})-\varepsilon)$ (56)
$=R(\bar{s},a^{*})+\gamma\rho_{\bar{s},a^{*}}(V^{*}(s^{\prime}))-\gamma\varepsilon\,$, (57)
@@
$R(\bar{s},a^{G})+\gamma\rho_{\bar{s},a^{G}}(\widehat{V}(s^{\prime}))\leq R(\bar{s},a^{G})+\gamma\rho_{\bar{s},a^{G}}(V^{*}(s^{\prime}))+\gamma\varepsilon,$ (58)
which collectively implie
$R(\bar{s},a^{*})-R(\bar{s},a^{G})\leq 2\gamma\varepsilon+\gamma\big(\rho_{\bar{s},a^{G}}(V^{*}(s^{\prime}))-\rho_{\bar{s},a^{*}}(V^{*}(s^{\prime})\big)\,.$ (59)

### v3

@@ D.1 Proof of Lemma 1
$V(s):=\max_{a}Q(s,a)\,,\qquad\ X(s):=\max_{a}W(s,a)\,.$
Assuming $\beta>0$, we then have
$\|Q^{\prime}-W^{\prime}\|$ $=Q^{\prime}(s,a)-W^{\prime}(s,a)$
$=\gamma\|V-X\|$
$\leq\gamma\|Q-W\|\,$,
and the lemma follows. The proof for the case of $\beta<0$ follows very similar lines, and is thus omitted. ∎
D.2 Proof of Lemma 2
Proof.
Solving $\frac{\gamma^{k}}{1-\gamma}$ for $k$, we get that if $k>\frac{-\log((1-\gamma)\varepsilon)}{\log(1/\gamma)}$, then $\|Q_{k}-Q^{*}\|<\varepsilon$, thus proving the first claim.
To show the other claim, we start by noting that $\|V^{\pi_{k}}-V^{*}\|\leq\|Q^{\pi_{k}}-Q^{*}\|$. Furthermore, by design we have that $\mathcal{T}^{\pi_{k}}Q^{\pi_{k}}=Q^{\pi_{k}}$ and that $\mathcal{T}Q_{k}=\mathcal{T}^{\pi_{k}}Q_{k}$. Thus,
$\|Q^{\pi_{k}}-Q^{*}\|\leq\|Q^{\pi_{k}}-Q_{k}\|+\|Q_{k}-Q^{*}\|\,.$
The first term in the right-hand side is bounded as follows:
$\|Q^{\pi_{k}}-Q_{k}\|$ $=\|\mathcal{T}^{\pi_{k}}Q^{\pi_{k}}-Q_{k}\|$
$\leq\gamma\|Q^{\pi_{k}}-Q_{k}\|+\gamma\|Q_{k}-Q_{k-1}\|\,$,
which means that
$\|Q^{\pi_{k}}-Q_{k}\|\leq\frac{\gamma}{1-\gamma}\|Q_{k}-Q_{k-1}\|\leq\frac{\gamma^{k}}{1-\gamma}\|Q_{1}-Q_{0}\|\leq\frac{\gamma^{k}}{(1-\gamma)^{2}}\,.$
The proof is completed by observing that picking $k>\log\big(\frac{2}{(1-\gamma)^{2}\varepsilon}\big)/\log(1/\gamma)$ implies $\|V^{\pi_{k}}-V^{*}\|<\varepsilon$. ∎
D.3 Proof of Lemma 3
Proof.
There are four cases to consider, corresponding to the combinations of $\beta>0$ or $\beta<0$ and whether, at the state–action pair $(s,a)$ attaining the maximum, $Q_{1}(s,a)>Q_{2}(s,a)$ or $Q_{2}(s,a)>Q_{1}(s,a)$.
Case 1: $\beta>0$ and $Q_{1}(s,a)>Q_{2}(s,a)$.
We have
$\|Q_{1}-Q_{2}\|$ $=\frac{\gamma}{\beta}\log\bigg(\frac{\sum_{s^{\prime}}P_{2}(s^{\prime}|s,a)e^{-\beta V_{2}(s^{\prime})}}{\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})}}\bigg)$
$=\frac{\gamma}{\beta}\log\bigg(\frac{\sum_{s^{\prime}}P_{2}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})+\beta(V_{2}(s^{\prime})-V_{1}(s^{\prime}))}}{\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})}}\bigg)$
$\leq\frac{\gamma}{\beta}\log\bigg(e^{\beta\|V_{1}-V_{2}\|}\frac{\sum_{s^{\prime}}P_{2}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})}}{\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})}}\bigg)$
$=\gamma\|V_{1}-V_{2}\|+\frac{\gamma}{\beta}\log\bigg(\frac{\sum_{s^{\prime}}P_{2}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})}}{\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})}}\bigg)$
$\leq\gamma\|Q_{1}-Q_{2}\|+\frac{\gamma}{\beta}\log\bigg(1+\frac{\sum_{s^{\prime}}P_{2}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})}-\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})}}{\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})}}\bigg)$
$\leq\gamma\|Q_{1}-Q_{2}\|+\frac{\gamma}{\beta}\frac{|\sum_{s^{\prime}}[P_{2}(s^{\prime}|s,a)-P_{1}(s^{\prime}|s,a)]e^{-\beta V_{1}(s^{\prime})}|}{e^{-\beta/(1-\gamma)}}\,.$
Rearranging the terms yields the asserted result:
$\|Q_{1}-Q_{2}\|\leq\frac{\gamma}{1-\gamma}\frac{e^{\beta/(1-\gamma)}}{\beta}\big|\sum_{s^{\prime}}[P_{2}(s^{\prime}|s,a)-P_{1}(s^{\prime}|s,a)]e^{-\beta V_{1}(s^{\prime})}\big|\,.$
Case 2: $\beta>0$ and $Q_{1}(s,a)<Q_{2}(s,a)$.
The proof is very similar to Case 1, but the extension $V_{2}(s)=V_{1}(s)+V_{2}(s)-V_{1}(s)$ is now done in the numerator instead:
$\|Q_{1}-Q_{2}\|$ $=\frac{\gamma}{\beta}\log\bigg(\frac{\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})}}{\sum_{s^{\prime}}P_{2}(s^{\prime}|s,a)e^{-\beta V_{2}(s^{\prime})}}\bigg)$
$=\frac{\gamma}{\beta}\log\bigg(\frac{\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})}}{\sum_{s^{\prime}}P_{2}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})-\beta(V_{2}(s^{\prime})-V_{1}(s^{\prime}))}}\bigg)$
$\leq\frac{\gamma}{\beta}\log\bigg(e^{\beta\|V_{1}-V_{2}\|}\frac{\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})}}{\sum_{s^{\prime}}P_{2}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})}}\bigg)$
$=\gamma\|V_{1}-V_{2}\|+\frac{\gamma}{\beta}\log\bigg(\frac{\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})}}{\sum_{s^{\prime}}P_{2}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})}}\bigg)$
$\leq\gamma\|Q_{1}-Q_{2}\|+\frac{\gamma}{\beta}\log\bigg(1+\frac{\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})}-\sum_{s^{\prime}}P_{2}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})}}{\sum_{s^{\prime}}P_{2}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})}}\bigg)$
$\leq\gamma\|Q_{1}-Q_{2}\|+\frac{\gamma}{\beta}\frac{|\sum_{s^{\prime}}[P_{1}(s^{\prime}|s,a)-P_{2}(s^{\prime}|s,a)]e^{-\beta V_{1}(s^{\prime})}|}{e^{-\beta/(1-\gamma)}},$
which again yields
$\|Q_{1}-Q_{2}\|\leq\frac{\gamma}{1-\gamma}\frac{e^{\beta/(1-\gamma)}}{\beta}\big|\sum_{s^{\prime}}[P_{2}(s^{\prime}|s,a)-P_{1}(s^{\prime}|s,a)]e^{-\beta V_{1}(s^{\prime})}\big|\,.$
Case 3: $\beta<0$ and $Q_{1}(s,a)>Q_{2}(s,a)$.
We have
$\|Q_{1}-Q_{2}\|$ $=\frac{\gamma}{|\beta|}\log\bigg(\frac{\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{|\beta|V_{1}(s^{\prime})}}{\sum_{s^{\prime}}P_{2}(s^{\prime}|s,a)e^{|\beta|V_{2}(s^{\prime})}}\bigg)$
$=\frac{\gamma}{|\beta|}\log\bigg(\frac{\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{|\beta|V_{1}(s^{\prime})}}{\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{|\beta|V_{1}(s^{\prime})-|\beta|(V_{2}(s^{\prime})-V_{1}(s^{\prime}))}}\bigg)$
$\leq\frac{\gamma}{|\beta|}\log\bigg(\frac{\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{|\beta|V_{1}(s^{\prime})}}{\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{|\beta|V_{1}(s^{\prime})-|\beta|\|V_{1}-V_{2}\|}}\bigg)$
$=\gamma\|V_{1}-V_{2}\|+\frac{\gamma}{|\beta|}\log\bigg(\frac{\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{|\beta|V_{1}(s^{\prime})}}{\sum_{s^{\prime}}P_{2}(s^{\prime}|s,a)e^{|\beta|V_{1}(s^{\prime})}}\bigg)$
$\leq\gamma\|Q_{1}-Q_{2}\|+\frac{\gamma}{|\beta|}\log\bigg(1+\frac{\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{|\beta|V_{1}(s^{\prime})}-\sum_{s^{\prime}}P_{2}(s^{\prime}|s,a)e^{|\beta|V_{1}(s^{\prime})}}{\sum_{s^{\prime}}P_{2}(s^{\prime}|s,a)e^{|\beta|V_{1}(s^{\prime})}}\bigg)$
$\leq\gamma\|Q_{1}-Q_{2}\|+\frac{\gamma}{|\beta|}e^{|\beta|/(1-\gamma)}\big|\sum_{s^{\prime}}[P_{2}(s^{\prime}|s,a)-P_{1}(s^{\prime}|s,a)]e^{|\beta|\big[V_{1}(s^{\prime})-\frac{1}{1-\gamma}\big]}\big|,$
which implies
$\|Q_{1}-Q_{2}\|\leq\frac{\gamma}{1-\gamma}\frac{e^{|\beta|/(1-\gamma)}}{|\beta|}\big|\sum_{s^{\prime}}[P_{2}(s^{\prime}|s,a)-P_{1}(s^{\prime}|s,a)]e^{|\beta|[V_{1}(s^{\prime})-\frac{1}{1-\gamma}]}\big|.$
Case 4: $\beta<0$ and $Q_{2}(s,a)>Q_{1}(s,a)$.
The proof of this case is similar to the other three cases and is omitted. ∎
D.4 Proof of Lemma 6.1
Let $N$ denote the number of calls to the generative model on each state-action pair such that the total number of calls is $SAN$. Let $\widehat{P}(s^{\prime}|s,a)$ denote the plug-in estimator obtained from $N$ samples of $s^{\prime}\sim P_{s,a}$, that is $\widehat{P}(s^{\prime}|s,a)=\frac{1}{N}\sum_{n=1}^{N}\mathbbmss{1}_{\{X_{n}=s^{\prime}\}}$, where $X_{n}$ taking values in $\mathcal{S}$ according to $P_{s,a}$.
@@ D.3 Proof of Lemma 3
We only prove the first claim —i.e., the case of $\beta>0$— as the other case is proven using completely similar lines. We note that for the random variable $\sum_{s^{\prime}}\mathbbmss{1}_{\{X_{n}=s^{\prime}\}}e^{-\beta V^{\pi}(s^{\prime})}$, we have that
$\mathbb{E}\bigg[\sum_{s^{\prime}}\mathbbmss{1}_{\{X_{n}=s^{\prime}\}}e^{-\beta V^{\pi}(s^{\prime})}\bigg]$ $=\sum_{s^{\prime}}\mathbb{E}\big[\mathbbmss{1}_{\{X_{n}=s^{\prime}\}}\big]e^{-\beta V^{\pi}(s^{\prime})}$
$=\sum_{s^{\prime}}P(s^{\prime}|s,a)e^{-\beta V^{\pi}(s^{\prime})}$
and that it is bounded in $e^{-\beta/(1-\gamma)},1]$. Also, since
$\sum_{s^{\prime}}\widehat{P}(s^{\prime}|s,a)e^{-\beta V^{\pi}(s^{\prime})}=\frac{1}{N}\sum_{n=1}^{N}\sum_{s^{\prime}}\mathbbmss{1}_{\{X_{n}=s^{\prime}\}}e^{-\beta V^{\pi}(s^{\prime})},$
it follows directly from Hoeffding’s inequality that
$\mathbb{P}\bigg(\bigg|\sum_{s^{\prime}}[P(s^{\prime}|s,a)-\widehat{P}(s^{\prime}|s,a)]e^{-\beta V^{\pi}(s^{\prime})}\bigg|\geq\varepsilon\bigg)\leq 2\exp\bigg(-\frac{2N\varepsilon^{2}}{\big(1-e^{-\beta/(1-\gamma)}\big)^{2}}\bigg).$
Thus, by picking $N=\frac{1}{2\varepsilon^{2}}\big(1-e^{-\beta/(1-\gamma)}\big)^{2}\log(2SA/\delta)$ and a union bound,
$\mathbb{P}\bigg(\max_{s,a}|\sum_{s^{\prime}}[P(s^{\prime}|s,a)-\widehat{P}(s^{\prime}|s,a)]e^{-\beta V^{\pi}(s^{\prime})}|\geq\varepsilon\bigg)\leq\delta\,.$
Appendix E Proof of Lemma 6.1.1
Next we prove a result that bounds the quality of a greedy policy with respect to the quality of the value-function for which the policy is greedy. The result is a generalization of [68] from the expectation to that of ERM and the derivation follow the same lines. Throughout, we use the notation $\rho_{s,a}(V(s^{\prime}))$ as shorthand notation for $\rho$ applied to the categorical random variable $X$ with support $\{V(s^{\prime})\}_{s^{\prime}\in\mathcal{S}}$ where $\mathbb{P}(X=V(s^{\prime}))=P(s^{\prime}|s,a)$.
[…]
Proof.
Let $\bar{s}$ be a state such that $\|V^{*}-V^{G}\|=V^{*}(\bar{s})-V^{G}(\bar{s})$, where $V^{G}:=V^{\pi_{G}}$. We then consider the two actions $a^{*}:=\pi^{*}(\bar{s})$ and $a^{G}:=\pi^{G}(\bar{s})$; ties can be breaking arbitrarily. Since $\pi^{G}$ is greedy with respect to $V^{G}$, we have that
$R(\bar{s},a^{*})+\gamma\rho_{\bar{s},a^{*}}(\widehat{V}(s^{\prime}))\leq R(\bar{s},a^{G})+\gamma\rho_{\bar{s},a^{G}}(\widehat{V}(s^{\prime}))\,.$
@@
$R(\bar{s},a^{*})+\gamma\rho_{\bar{s},a^{*}}(\widehat{V}(s^{\prime}))$ $\geq R(\bar{s},a^{*})+\gamma\rho_{\bar{s},a^{*}}(V^{*}(s^{\prime})-\varepsilon)$
$=R(\bar{s},a^{*})+\gamma\rho_{\bar{s},a^{*}}(V^{*}(s^{\prime}))-\gamma\varepsilon\,$,
@@
$R(\bar{s},a^{G})+\gamma\rho_{\bar{s},a^{G}}(\widehat{V}(s^{\prime}))\leq R(\bar{s},a^{G})+\gamma\rho_{\bar{s},a^{G}}(V^{*}(s^{\prime}))+\gamma\varepsilon,$
which collectively imply
$R(\bar{s},a^{*})-R(\bar{s},a^{G})\leq 2\gamma\varepsilon+\gamma\big(\rho_{\bar{s},a^{G}}(V^{*}(s^{\prime}))-\rho_{\bar{s},a^{*}}(V^{*}(s^{\prime})\big)\,.$